# NoMIRACL: Knowing When You Don’t Know for Robust Multilingual Retrieval-Augmented Generation

### v2

@@ Acknowledgements
We would like to thank our annotators, without whom NoMIRACL could not have been built. We would also like to thank Akintunde Oladipo for providing necessary Microsoft Azure credits for evaluating OpenAI models. This research was supported in part by the Natural Sciences and Engineering Research Council (NSERC) of Canada, a gift from Huawei, and Cloud TPU support from Google’s TPU Research Cloud (TRC).
@@ Appendix D Quality control
Phase I. In this phase, reviewers were given both the prompts and the generated queries and filled up a checklist to determine whether the quality of the queries met our requirements. Criteria include the examination of the query itself (e.g., spelling, syntax, and fluency, etc.) and whether the query could be answered directly by the prompt, which we wanted to avoid to generate more informative queries, following Clark et al. (2020); Zhang et al. (2023). In order to evaluate this, we measured the lexical overlap between the queries and their corresponding prompts. We found the overlaps primarily occur in entities or stopwords and thus concluded that the generated queries are reasonably different from the given prompts.
@@
Appendix E Model Checkpoints and Results
All the LLM checkpoints used in our experiments, for both closed and open-sourced, can be found in Table 3. Additionally, LLM evaluation results for both the non-relevant and relevant subset across all languages can be found in Figure 8 and Figure 9 respectively.
Table 3: LLM checkpoint links used for evaluation in NoMIRACL.
@@ Appendix E Model Checkpoints and Results
Figure 8: Hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}$) on the non-relevant subset in NoMIRACL test split. The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ (where $k=10$) passages are judged by a human annotator as non-relevant. On average, a majority of LLMs (except Mistral) hallucinate on the non-relevant subset. Lower the hallucination rate is better.
@@
Figure 9: Error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$) on the relevant subset in NoMIRACL test split. The relevant subset contains queries with known answers, i.e., at least one of the top-$k$ (where $k=10$) passages judged by a human annotator is relevant. On average, a majority of LLMs (except Mistral and Aya-101) have a lower error rate, i.e., can accurately identify the relevant answer. Lower the error rate is better.
Language Query Judged Passages Relevance
English (en)
What is the name of the winner of the Abel Prize of 2022?
@@
[1] Abel Prize: The Abel Prize () is a Norwegian prize awarded annually by the King of Norway to one or more outstanding mathematicians. It is named after Norwegian mathematician Niels Henrik Abel (1802–1829) and directly modeled after the Nobel Prizes. It comes with a monetary award of 6 million Norwegian kroner (NOK) (€635,000 or \$740,000).
0
@@
[2] MIT Department of Mathematics: The current faculty of around 50 members includes Wolf Prize winner Michael Artin, Abel Prize winner Isadore Singer, and numerical analyst Gilbert Strang.
0
@@
[3] Abel Prize: After interest in the concept of the prize had risen in 2001, a working group was formed to develop a proposal, which was presented to the Prime Minister of Norway in May. In August 2001, the Norwegian government announced that the prize would be awarded beginning in 2002, the two-hundredth anniversary of Abel’s birth. Atle Selberg received an honorary Abel Prize in 2002, but the first actual Abel Prize was awarded in 2003.
0
@@
[4] Abel Prize: The prize was first proposed in 1899, to be part of the celebration of the 100th anniversary of Niels Henrik Abel’s birth in 1902. Shortly before his death in 1899, the Norwegian mathematician Sophus Lie proposed establishing an Abel Prize when he learned that Alfred Nobel’s plans for annual prizes would not include a prize in mathematics. King Oscar II was willing to finance a mathematics prize in 1902, and the mathematicians Ludwig Sylow and Carl Størmer drew up statutes and rules for the proposed prize. However, Lie’s influence waned after his death, and the dissolution of the union between Sweden and Norway in 1905 ended the first attempt to create an Abel Prize.
0
@@
[5] Eötvös Loránd University: Eötvös Loránd University (, ELTE) is a Hungarian public research university based in Budapest. Founded in 1635, ELTE is one of the largest and most prestigious public higher education institutions in Hungary. The 28,000 students at ELTE are organized into eight faculties, and into research institutes located throughout Budapest and on the scenic banks of the Danube. ELTE is affiliated with 5 Nobel laureates, as well as winners of the Wolf Prize, Fulkerson Prize and Abel Prize, the latest of which was Abel Prize winner Endre Szemerédi in 2012.
0
@@
[6] Abel Prize: Anyone may submit a nomination for the Abel Prize, however, self-nominations are not permitted. The nominee must be alive; however, if the awardee dies after being declared as the winner, the prize will be awarded posthumously.
0
@@
[7] Abel Prize: The Norwegian Academy of Science and Letters declares the winner of the Abel Prize each March after recommendation by the Abel Committee, which consists of five leading mathematicians. Both Norwegians and non-Norwegians may serve on the Committee. They are elected by the Norwegian Academy of Science and Letters and nominated by the International Mathematical Union and the European Mathematical Society. The committee is of 2018 chaired by Norwegian mathematician Hans Munthe-Kaas (University of Bergen), and was before that, headed by Professor John Rognes.
0
@@
[8] Hans Munthe-Kaas: Munthe-Kaas received Exxon Mobil Award for best PhD at NTNU, 1989, and the Carl-Erik Frōberg Prize in Numerical Mathematics 1996 for the paper L̈ie–Butcher theory for Runge–Kutta Methods.̈ Munthe-Kaas is elected member of the Norwegian Academy of Science and Letters, the Royal Norwegian Society of Sciences and Letters and the Norwegian Academy of Technological Sciences. Munthe-Kaas is the chair of the international Abel prize committee (2018-2022), he is President of the Scientific Council of Centre International de Mathématiques Pures et Appliquées (CIMPA) (2017–present) and he is Editor-in-Chief of Journal Foundations of Computational Mathematics (2017–present). Munthe-Kaas was secretary of Foundations of Computational Mathematics (2005–2011) and member of the Board of the Abel Prize in Mathematics (2010–2018).
0
@@
[9] Science and technology in Hungary: Among Hungary’s numerous research universities, the Eötvös Loránd University, founded in 1635, is one of the largest and the most prestigious public higher education institutions in Hungary. The 28,000 students at ELTE are organized into eight faculties, and into research institutes located throughout Budapest. ELTE is affiliated with 5 Nobel laureates, as well as winners of the Wolf Prize, Fulkerson Prize and Abel Prize, the latest of which was Abel Prize winner Endre Szemerédi in 2012.
0
@@
[10] Abel Prize: The Abel Prize’s history dates back to 1899, when its establishment was proposed by the Norwegian mathematician Sophus Lie when he learned that Alfred Nobel’s plans for annual prizes would not include a prize in mathematics. In 1902 King Oscar II of Sweden and Norway indicated his willingness to finance a mathematics prize to complement the Nobel Prizes, but the establishment of the prize was prevented by the dissolution of the union between Norway and Sweden in 1905. It took almost a century before the prize was finally established by the Government of Norway in 2001, and it was specifically intended ẗo give the mathematicians their own equivalent of a Nobel Prize.T̈he laureates are selected by the Abel Committee, the members of which are appointed by the Norwegian Academy of Science and Letters.
0
@@
Table 4: Randomly sampled example of a query and 10 judged passages in English (en) from the non-relevant subset (test split) in NoMIRACL. Titles of each passage are marked in bold. The relevance judgement has been annotated by a native speaker manually.
Language Query Judged Passages Relevance
English (en)
In which country Praia dos Pescadores is?
@@
[1] Praia dos Pescadores (Albufeira): Praia dos Pescadores or the “Fishermans Beach” is a blue flag beach on the Atlantic south coast of the Algarve, in the district of Bairro dos Pescadores (Neighborhood of the Fisherman), Albufeira which is within the Municipality of Albufeira, Portugal. The beach is one of the two beaches which front the town of “Albufeira” with “Praia do Túnel” at the western end and “Praia dos Pescadores"" lying to the eastern end of the towns seafront. The town and its beaches are located west by road of the regions capital of Faro. In the days before Albufeira had a harbour and mariner the “Praia dos Pescadores” was where all the local fishermen operated from and the beach scene would have been very different to the site you see today. Then the beach would have been full of brightly painted fishing boats pulled up on this beach when not at sea and much of the tourist activities took place on the “Praia do Túnel”. Today the “Praia dos Pescadores” is now used for tourism and is a very busy beach especially in the summer season.
1
@@
[2] Praia do Túnel (Peneco): Praia do Túnel is a beach on the Atlantic south coast of the Algarve, in the town of Albufeira which is within the Municipality of Albufeira, Portugal. The beach is also known as “Praia do Peneco” and is one of the two beaches which front the town of “Albufeira” with “Praia do Túnel” at the western end and “Praia dos Pescadores” lying to the eastern end of the towns seafront. The town and its beaches are located west by road of the regions capital of Faro. The beach gets its name from a 20 meter long tunnel next to the tourist office in the middle of Albufeira which cuts through the cliffs linking the towns square to the beach. At the western end of the beach there is a promenade which ends at the cave known as the Xorino Grotto. According to 13th-century legend, the cave was used as shelter by the Moors after the Christian conquest of Albufeira. As well as the tunnel there are several other points of access to the beach including a lift, ramps and steps […]
0
@@
[3] Praia dos Pescadores (Albufeira): The beach is in length and is wide at low tide. The beach is divided by a protruding cliff from Praia do Túnel at the western end of the seafront. To the beaches eastern boundary is the Praia do Inatel and it is divided from that beach by a concrete pier which covers the outflow of the Ribeira de Albufeira (Albufeira River). There are also cliffs at the eastern end and to the back of the beach there is an amphitheatre of white houses of the district of Bairro dos Pescadores. The beach can also be accessed by an outdoor foot escalator from the Pau da Bandeira bluff located south of Bairro dos Pescadores down to the beach and Albufeira old town.
0
@@
[4] Praia dos Pescadores (Albufeira): Praia dos Pescadores is an easily accessed beach with its large hard surface square at beach level. There are two car park’s near-by, one of which, is at beach level, a short distance along the Avenida 25 de Abril within the old town. The second car park is at the top of the cliffs at Bairro dos Pescadores and is accessed via the outdoor escalator. To the back of the western end of the beach there a variety of restaurants many of which specialise in the local fish and seafood. The beach has several licensed concessions with opportunities to hire parasols and sun loungers. There are also many organised beach and water sport concessions from volleyball to boat trips and Parasailing. The beach also has toilet and shower facilities. During the summer months the beach is patrolled by lifeguards. In recent years the beach has been the focal point for the new year celebrations in the town. A temporary concert stage is erected on the Largo 25 de Abril and concerts are held to celebrate the new year. In the past the celebration has seen international bands appearing such as British reggae/pop band UB40 in 2009. The celebrations cumulate with a firework display held just of the beach on boats and pontoons just of the shoreline.
0
@@
[5] Praia do Penedo: Praia do Penedo is a beach within the Municipality of Aljezur, in the Algarve, Portugal. The beach is on the western Seaboard in the north west of the Algarve. The beach is south west of the village of Aljezur, and is north west, by road, from the regions capital of Faro. The beach of Praia do Penedo is inside the Vicentine Coast Natural Park, an area of outstanding natural beauty.
0
@@
[6] Praia do Norte: Praia do Norte is a civil parish of the municipality of Horta, located along the northern coast between Cedros and Capelo, on the Portuguese island of Faial, in the archipelago of the Azores. The population in 2011 was 250, in an area of . It is the least populous parish on the island, reached along the Ëstrada RegionalË.R.1-1ª regional roadway from the urban centre of Horta.
0
@@
[7] Praia das Conchas, São Tomé and Príncipe: Praia das Conchas is a settlement in the western part of the Lobata District on São Tomé Island in São Tomé and Príncipe. Its population is 174 (2012 census). Established as a plantation (r̈oça)̈, Praia das Conchas lies 2 km from the coast, 3 km west of Guadalupe. There is a smaller seaside settlement also called P̈raia das Conchas,̈ 3.5 km to the north.
0
@@
[8] Praia das Gatas: Praia das Gatas (Portuguese meaning b̈each of the cats)̈ is a sandy beach in the northeastern part of the island of Boa Vista in Cape Verde. The nearest village is Fundo das Figueiras, 5km to the southwest. It forms a part of Northern Nature Park (P̈arque Natural do Norte)̈. The small island Ilhéu dos Pássaros lies off the coast at the Praia das Gatas.
0
@@
[9] Praia (Santa Cruz da Graciosa): Praia (officially São Mateus da Praia) is a Portuguese civil parish in the municipality of Santa Cruz da Graciosa, on the island of Graciosa, in the Azores. It still retains its former name locally, owing to the parish having once been the historical municipality of Praia. The population in 2011 was 836, in an area of 12.82 km².
0
@@
[10] Praia Harbor: Praia Harbor () is the port of the city of Praia in the southern part of the island of Santiago, Cape Verde. It is situated in a natural bay of the Atlantic Ocean. Since the latest modernization in 2014, it has 2 long quays, 3 shorter quays, a quay for fishing boats with fish processing installations, 2 container parks, 2 roll-on/roll-off ramps and a passenger terminal. The total length of the quays is 863 m, and the maximum depth is 13.5 m. The port of Praia played an important role in the colonization of Africa and South America by the Portuguese. With 817,845 metric tonnes of cargo and 85,518 passengers handled (2017), it is the second busiest port of Cape Verde, after Porto Grande (Mindelo).
0
@@
Table 5: Randomly sampled example of a query and 10 judged passages in English (en) from the relevant subset (test split) in NoMIRACL. Titles of each passage are marked in bold. The relevance judgement has been annotated by a native speaker manually.

### v3

@@ Acknowledgements
We would like to thank our annotators, without whom NoMIRACL could not have been built. We would also like to thank Akintunde Oladipo for providing the necessary Microsoft Azure credits for evaluating OpenAI models. This research was supported in part by the Natural Sciences and Engineering Research Council (NSERC) of Canada, a gift from Huawei, and Cloud TPU support from Google’s TPU Research Cloud (TRC).
@@ Appendix D Quality Control
Phase I. In this phase, reviewers were given both the prompts and the generated queries and filled up a checklist to determine whether the quality of the queries met our requirements. Criteria include the examination of the query itself (e.g., spelling, syntax, and fluency, etc.) and whether the query could be answered directly by the prompt, which we wanted to avoid to generate more informative queries, following Clark et al. (2020); Zhang et al. (2023). To evaluate this, we measured the lexical overlap between the queries and their corresponding prompts. We found the overlaps primarily occur in entities or stopwords and thus concluded that the generated queries are reasonably different from the given prompts.
@@
Appendix E Checkpoints and Additional Results
All multilingual-focused LLM checkpoints used in our experiments for both closed and open-sourced can be found in Appendix E. Hyperparameter choices during NoMIRACL supervised fine-tuning LLMs are listed in Table 6 and experimental results in Table 7. LLM evaluation results for both the non-relevant and relevant subsets for all models can be found in Figure 8 and Figure 9 respectively. section 6 shows template changes for prompt optimization ablation experiments, including (i) role, (ii) repeat, and (iii) explanation prompts.
Table 5: All models and checkpoint links used for NoMIRACL evaluation.
Hyperparameter Value
use_peft true
torch_dtype bfloat16
lora_r 64
lora_alpha 16
lora_dropout 0.05
lora_target_modules {q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj}
learning_rate 3.0e-06
lr_scheduler_type cosine
max_seq_length 4096
Table 6: Hyperparameter settings chosen during LoRA supervised fine-tuning (SFT) Mistral-7B (v0.3) and LLAMA-3 (8B) instruct models on the NoMIRACL development split.
@@ Appendix E Checkpoints and Additional Results
Figure 8: Hallucination rate (in %) = $\mathrm{FP}/(\mathrm{FP}+\mathrm{TN}$) on the non-relevant subset in NoMIRACL test split. The non-relevant subset contains queries with no-known answers, i.e., all top-$k$ (where $k=10$) passages are judged by a human annotator as non-relevant. On average, most LLMs (except Mistral) hallucinate on the non-relevant subset. Lower the hallucination rate is better.
@@
Figure 9: Error rate (in %) = $\mathrm{FN}/(\mathrm{FN}+\mathrm{TP}$) on the relevant subset in NoMIRACL test split. The relevant subset contains queries with known answers, i.e., at least one of the top-$k$ (where $k=10$) passages judged by a human annotator is relevant. On average, most LLMs (except Mistral and Aya-101) have a lower error rate, i.e., can accurately identify the relevant answer. Lower the error rate is better.
QUESTION: {query} CONTEXTS: [1] {Passage title}: {Passage text} [2] {Passage title}: {Passage text} ... [10] {Passage title}: {Passage text} Please remember to read all the contexts carefully. If any of the contexts answers the question: {query}, respond as either ‘‘Yes, answer is present’’ or ‘‘I don’t know’’. OUTPUT:
{mdframed}
[backgroundcolor=gray!5] Read the query and the contexts carefully and provide a step-by-step explanation for your answer. If any of the contexts answers the question, respond as either ‘‘Yes, answer is present’’ or ‘‘I don’t know’’. You must strictly follow the output format with ## Reasoning: ... ## Answer: ‘‘Yes, answer is present’’ OR ‘‘I don’t know’’. QUESTION: {query} CONTEXTS: [1] {Passage title}: {Passage text} [2] {Passage title}: {Passage text} ... [10] {Passage title}: {Passage text} OUTPUT:
Figure 10: All prompt ablations used in our experiments for LLM hallucination evaluation for all 18 languages in NoMIRACL on both the relevant and non-relevant subsets. Role prompt appends the role of the LLM at the beginning of the prompt (highlighted in blue). Repeat prompt highlights the task by repeating instructions at the end of the prompt (highlighted in red). Explanation prompt asks the model to provide a reasoning path and finally answer the question (highlighted in violet).
ar bn de en es fa fr fi hi id ja ko ru sw te th yo zh Avg.
Hallucination Rates (in %) on NoMIRACL test split (non-relevant subset)
Llama-3 (8B) 19.6 20.0 29.5 19.2 30.0 44.8 35.2 10.5 14.4 23.2 52.0 31.6 26.8 4.0 13.2 41.6 1.6 64.4 26.8
Llama-3 (8B) (w/ SFT) 83.0 12.5 36.9 44.0 46.8 65.6 36.9 70.0 61.5 53.1 41.7 9.7 85.4 42.2 0.0 82.7 12.0 20.0 44.7
Mistral-7B (v0.3) 40.4 63.2 38.2 42.8 17.2 52.4 47.6 16.1 39.6 30.8 44.4 28.8 41.6 14.8 74.0 58.8 23.6 45.2 40.0
Mistral-7B (v0.3) (w/ SFT) 46.8 33.2 34.1 73.6 33.6 52.4 48.0 42.7 26.0 59.2 52.0 47.6 62.8 35.6 31.2 40.0 45.2 33.2 44.3
Error Rates (in %) on NoMIRACL test split (relevant subset)
Llama-3 (8B) 53.6 56.0 36.8 41.6 26.8 26.8 34.4 61.2 56.8 48.4 17.6 35.6 46.4 80.8 65.6 29.2 91.7 6.0 45.3
Llama-3 (8B) (w/ SFT) 6.0 70.8 32.4 31.2 24.0 14.6 21.2 8.8 14.9 40.6 24.0 77.2 8.3 32.5 87.5 4.2 75.5 50.4 34.7
Mistral-7B (v0.3) 14.4 20.4 21.6 8.0 28.0 22.4 15.6 38.0 30.8 43.6 16.0 30.0 18.0 57.6 17.6 13.6 50.0 13.2 25.5
Mistral-7B (v0.3) (w/ SFT) 46.4 60.8 46.0 12.4 46.8 36.4 35.6 42.0 69.6 37.2 37.6 42.0 25.2 69.6 66.4 56.8 52.0 46.8 46.1
Table 7: Complete SFT results using the NoMIRACL development dataset for LLAMA-3 (8B) and Mistral-7B (v0.3) LLMs across all languages in NoMIRACL. Lower the hallucination and error rates (%) is better.
Query Judged Passages Relevance
@@
What is the name of the winner of the Abel Prize of 2022? [1] Abel Prize: The Abel Prize () is a Norwegian prize awarded annually by the King of Norway to one or more outstanding mathematicians. It is named after Norwegian mathematician Niels Henrik Abel (1802–1829) and directly modeled after the Nobel Prizes. It comes with a monetary award of 6 million Norwegian kroner (NOK) (€635,000 or \$740,000). 0
@@
[2] MIT Department of Mathematics: The current faculty of around 50 members includes Wolf Prize winner Michael Artin, Abel Prize winner Isadore Singer, and numerical analyst Gilbert Strang. 0
@@
[3] Abel Prize: After interest in the concept of the prize had risen in 2001, a working group was formed to develop a proposal, which was presented to the Prime Minister of Norway in May. In August 2001, the Norwegian government announced that the prize would be awarded beginning in 2002, the two-hundredth anniversary of Abel’s birth. Atle Selberg received an honorary Abel Prize in 2002, but the first actual Abel Prize was awarded in 2003. 0
@@
[4] Abel Prize: The prize was first proposed in 1899, to be part of the celebration of the 100th anniversary of Niels Henrik Abel’s birth in 1902. Shortly before his death in 1899, the Norwegian mathematician Sophus Lie proposed establishing an Abel Prize when he learned that Alfred Nobel’s plans for annual prizes would not include a prize in mathematics. King Oscar II was willing to finance a mathematics prize in 1902, and the mathematicians Ludwig Sylow and Carl Størmer drew up statutes and rules for the proposed prize. However, Lie’s influence waned after his death, and the dissolution of the union between Sweden and Norway in 1905 ended the first attempt to create an Abel Prize. 0
@@
[5] Eötvös Loránd University: Eötvös Loránd University (, ELTE) is a Hungarian public research university based in Budapest. Founded in 1635, ELTE is one of the largest and most prestigious public higher education institutions in Hungary. The 28,000 students at ELTE are organized into eight faculties, and into research institutes located throughout Budapest and on the scenic banks of the Danube. ELTE is affiliated with 5 Nobel laureates, as well as winners of the Wolf Prize, Fulkerson Prize and Abel Prize, the latest of which was Abel Prize winner Endre Szemerédi in 2012. 0
@@
[6] Abel Prize: Anyone may submit a nomination for the Abel Prize, however, self-nominations are not permitted. The nominee must be alive; however, if the awardee dies after being declared as the winner, the prize will be awarded posthumously. 0
@@
[7] Abel Prize: The Norwegian Academy of Science and Letters declares the winner of the Abel Prize each March after recommendation by the Abel Committee, which consists of five leading mathematicians. Both Norwegians and non-Norwegians may serve on the Committee. They are elected by the Norwegian Academy of Science and Letters and nominated by the International Mathematical Union and the European Mathematical Society. The committee is of 2018 chaired by Norwegian mathematician Hans Munthe-Kaas (University of Bergen), and was before that, headed by Professor John Rognes. 0
@@
[8] Hans Munthe-Kaas: Munthe-Kaas received Exxon Mobil Award for best PhD at NTNU, 1989, and the Carl-Erik Frōberg Prize in Numerical Mathematics 1996 for the paper L̈ie–Butcher theory for Runge–Kutta Methods.̈ Munthe-Kaas is elected member of the Norwegian Academy of Science and Letters, the Royal Norwegian Society of Sciences and Letters and the Norwegian Academy of Technological Sciences. Munthe-Kaas is the chair of the international Abel prize committee (2018-2022), he is President of the Scientific Council of Centre International de Mathématiques Pures et Appliquées (CIMPA) (2017–present) and he is Editor-in-Chief of Journal Foundations of Computational Mathematics (2017–present). Munthe-Kaas was secretary of Foundations of Computational Mathematics (2005–2011) and member of the Board of the Abel Prize in Mathematics (2010–2018). 0
@@
[9] Science and technology in Hungary: Among Hungary’s numerous research universities, the Eötvös Loránd University, founded in 1635, is one of the largest and the most prestigious public higher education institutions in Hungary. The 28,000 students at ELTE are organized into eight faculties, and into research institutes located throughout Budapest. ELTE is affiliated with 5 Nobel laureates, as well as winners of the Wolf Prize, Fulkerson Prize and Abel Prize, the latest of which was Abel Prize winner Endre Szemerédi in 2012. 0
@@
[10] Abel Prize: The Abel Prize’s history dates back to 1899, when its establishment was proposed by the Norwegian mathematician Sophus Lie when he learned that Alfred Nobel’s plans for annual prizes would not include a prize in mathematics. In 1902 King Oscar II of Sweden and Norway indicated his willingness to finance a mathematics prize to complement the Nobel Prizes, but the establishment of the prize was prevented by the dissolution of the union between Norway and Sweden in 1905. It took almost a century before the prize was finally established by the Government of Norway in 2001, and it was specifically intended ẗo give the mathematicians their own equivalent of a Nobel Prize.T̈he laureates are selected by the Abel Committee, the members of which are appointed by the Norwegian Academy of Science and Letters. 0
@@
Table 8: Randomly sampled example of a query on “What is the name of the winner of the Abel Prize of 2022?” and top-10 judged passages in English (en) from the non-relevant subset (test split) in NoMIRACL. Titles of each passage are marked in bold. The relevance judgment has been annotated manually by a native speaker.
Query Judged Passages Relevance
@@
In which country Praia dos Pescadores is? [1] Praia dos Pescadores (Albufeira): Praia dos Pescadores or the “Fishermans Beach” is a blue flag beach on the Atlantic south coast of the Algarve, in the district of Bairro dos Pescadores (Neighborhood of the Fisherman), Albufeira which is within the Municipality of Albufeira, Portugal. The beach is one of the two beaches which front the town of “Albufeira” with “Praia do Túnel” at the western end and “Praia dos Pescadores"" lying to the eastern end of the towns seafront. The town and its beaches are located west by road of the regions capital of Faro. In the days before Albufeira had a harbour and mariner the “Praia dos Pescadores” was where all the local fishermen operated from and the beach scene would have been very different to the site you see today. Then the beach would have been full of brightly painted fishing boats pulled up on this beach when not at sea and much of the tourist activities took place on the “Praia do Túnel”. Today the “Praia dos Pescadores” is now used for tourism and is a very busy beach especially in the summer season. 1
@@
[2] Praia do Túnel (Peneco): Praia do Túnel is a beach on the Atlantic south coast of the Algarve, in the town of Albufeira which is within the Municipality of Albufeira, Portugal. The beach is also known as “Praia do Peneco” and is one of the two beaches which front the town of “Albufeira” with “Praia do Túnel” at the western end and “Praia dos Pescadores” lying to the eastern end of the towns seafront. The town and its beaches are located west by road of the regions capital of Faro. The beach gets its name from a 20 meter long tunnel next to the tourist office in the middle of Albufeira which cuts through the cliffs linking the towns square to the beach. At the western end of the beach there is a promenade which ends at the cave known as the Xorino Grotto. According to 13th-century legend, the cave was used as shelter by the Moors after the Christian conquest of Albufeira. As well as the tunnel there are several other points of access to the beach including a lift, ramps and steps […] 0
@@
[3] Praia dos Pescadores (Albufeira): The beach is in length and is wide at low tide. The beach is divided by a protruding cliff from Praia do Túnel at the western end of the seafront. To the beaches eastern boundary is the Praia do Inatel and it is divided from that beach by a concrete pier which covers the outflow of the Ribeira de Albufeira (Albufeira River). There are also cliffs at the eastern end and to the back of the beach there is an amphitheatre of white houses of the district of Bairro dos Pescadores. The beach can also be accessed by an outdoor foot escalator from the Pau da Bandeira bluff located south of Bairro dos Pescadores down to the beach and Albufeira old town. 0
@@
[4] Praia dos Pescadores (Albufeira): Praia dos Pescadores is an easily accessed beach with its large hard surface square at beach level. There are two car park’s near-by, one of which, is at beach level, a short distance along the Avenida 25 de Abril within the old town. The second car park is at the top of the cliffs at Bairro dos Pescadores and is accessed via the outdoor escalator. To the back of the western end of the beach there a variety of restaurants many of which specialise in the local fish and seafood. The beach has several licensed concessions with opportunities to hire parasols and sun loungers. There are also many organised beach and water sport concessions from volleyball to boat trips and Parasailing. The beach also has toilet and shower facilities. During the summer months the beach is patrolled by lifeguards. In recent years the beach has been the focal point for the new year celebrations in the town. A temporary concert stage is erected on the Largo 25 de Abril and concerts are held to celebrate the new year. In the past the celebration has seen international bands appearing such as British reggae/pop band UB40 in 2009. The celebrations cumulate with a firework display held just of the beach on boats and pontoons just of the shoreline. 0
@@
[5] Praia do Penedo: Praia do Penedo is a beach within the Municipality of Aljezur, in the Algarve, Portugal. The beach is on the western Seaboard in the north west of the Algarve. The beach is south west of the village of Aljezur, and is north west, by road, from the regions capital of Faro. The beach of Praia do Penedo is inside the Vicentine Coast Natural Park, an area of outstanding natural beauty. 0
@@
[6] Praia do Norte: Praia do Norte is a civil parish of the municipality of Horta, located along the northern coast between Cedros and Capelo, on the Portuguese island of Faial, in the archipelago of the Azores. The population in 2011 was 250, in an area of . It is the least populous parish on the island, reached along the Ëstrada RegionalË.R.1-1ª regional roadway from the urban centre of Horta. 0
@@
[7] Praia das Conchas, São Tomé and Príncipe: Praia das Conchas is a settlement in the western part of the Lobata District on São Tomé Island in São Tomé and Príncipe. Its population is 174 (2012 census). Established as a plantation (r̈oça)̈, Praia das Conchas lies 2 km from the coast, 3 km west of Guadalupe. There is a smaller seaside settlement also called P̈raia das Conchas,̈ 3.5 km to the north. 0
@@
[8] Praia das Gatas: Praia das Gatas (Portuguese meaning b̈each of the cats)̈ is a sandy beach in the northeastern part of the island of Boa Vista in Cape Verde. The nearest village is Fundo das Figueiras, 5km to the southwest. It forms a part of Northern Nature Park (P̈arque Natural do Norte)̈. The small island Ilhéu dos Pássaros lies off the coast at the Praia das Gatas. 0
@@
[9] Praia (Santa Cruz da Graciosa): Praia (officially São Mateus da Praia) is a Portuguese civil parish in the municipality of Santa Cruz da Graciosa, on the island of Graciosa, in the Azores. It still retains its former name locally, owing to the parish having once been the historical municipality of Praia. The population in 2011 was 836, in an area of 12.82 km². 0
@@
[10] Praia Harbor: Praia Harbor () is the port of the city of Praia in the southern part of the island of Santiago, Cape Verde. It is situated in a natural bay of the Atlantic Ocean. Since the latest modernization in 2014, it has 2 long quays, 3 shorter quays, a quay for fishing boats with fish processing installations, 2 container parks, 2 roll-on/roll-off ramps, and a passenger terminal. The total length of the quays is 863 m, and the maximum depth is 13.5 m. The port of Praia played an important role in the colonization of Africa and South America by the Portuguese. With 817,845 metric tonnes of cargo and 85,518 passengers handled (2017), it is the second busiest port of Cape Verde, after Porto Grande (Mindelo). 0
@@
Table 9: Randomly sampled example of a query on “In which country Praia dos Pescadores is?” and top-10 judged passages in English (en) from the relevant subset (test split) in NoMIRACL. Titles of each passage are marked in bold. The relevance judgment has been annotated manually by a native speaker.